# Fast Trac: A Parameter-Free Optimizer for Lifelong Reinforcement Learning

### v2

@@ 1 Introduction
Spot’s struggle exemplifies a critical challenge in real-world decision making, known as lifelong reinforcement learning (lifelong RL). The idea is that the learning agent must continually acquire new knowledge to handle the nonstationarity of the environment. At first glance, there appears to be an obvious solution: given a policy gradient oracle, the agent could just keep running gradient descent nonstop. However, recent experiments have demonstrated an intriguing behavior called loss of plasticity (Dohare et al., 2021; Lyle et al., 2022; Abbas et al., 2023; Sokar et al., 2023): despite persistent gradient steps, such an agent can gradually lose its responsiveness to incoming observations. There are even extreme cases of loss of plasticity (known as negative transfer or primacy bias), where prior learning can significantly hamper the performance in new tasks (Nikishin et al., 2022; Ahn et al., 2024); see Figure 1 for an example. All these suggest that the problem is more involved than one might think.
@@
  • Algorithm: Trac Building on a series of results in OCO (Cutkosky & Orabona, 2018; Cutkosky, 2019; Cutkosky et al., 2023; Zhang et al., 2024b), we propose a (hyper)-parameter-free optimizer for lifelong RL, called Trac (AdapTive RegularizAtion in Continual environments). Intuitively, the idea is a refinement of regularization: instead of manually selecting the magnitude of regularization beforehand, Trac chooses that in an online, data-dependent manner. From the perspective of OCO theory, Trac is insensitive to its own hyperparameter, which means that no hyperparameter tuning is necessary in practice. Furthermore, as an optimization approach to lifelong RL, Trac is compatible with any policy parameterization.
  • Experiment Using Proximal Policy Optimization (PPO), we conduct comprehensive experiments on the instantiation of Trac called Trac PPO. A diverse range of lifelong RL environments are tested (based on Procgen, Atari, and Gym Control), with considerably larger scale than prior works. In settings where existing approaches (Abbas et al., 2023; Kumar et al., 2023) struggle, we find that Trac PPO
@@ 2 Lifelong RL
Modern RL approaches, including PPO (Schulman et al., 2017) and others, crucially utilize the idea of policy parameterization. Specifically, a policy refers to the distribution of the agent’s action $a_{t}$ (conditioned on the historical observations), and we use $\theta_{t}\in\mathbb{R}^{d}$ to denote the parameterizing weight vector. After sampling $a_{t}$ and receiving new observations, the agent could define a loss function $J_{t}(\theta)$ that characterizes the “hypothetical performance” of each weight $\theta\in\mathbb{R}^{d}$. Then, by computing the policy gradient $g_{t}=\nabla J_{t}(\theta_{t})$, one could apply a first order optimization algorithm [^3] Formally, a dynamical system that given its state $\theta_{t}$ and input $g_{t}$ outputs the new state $\textsc{OPT}(\theta_{t},g_{t})$. OPT to obtain the updated weight, $\theta_{t+1}=\textsc{OPT}(\theta_{t},g_{t})$.
@@
In the RL literature, the use of “lifelong” and “continual” varies significantly across studies, which may lead to confusion. Abel et al. (2024) characterized continual reinforcement learning (CRL) as a never-ending learning process. However, much of the literature cited under CRL, such as (Abbas et al., 2023; Ahn et al., 2024), primarily focuses on the problem of backward transfer (avoiding catastrophic forgetting). Conversely, the present work addresses the problem of forward transfer, which refers to the rapid adaptation to new tasks. Because of this we use “lifelong” rather than “continual” in our exposition, similar to (Thrun, 1996; Abel et al., 2018b; Julian et al., 2020).
@@ 3 Method
Trac for lifelong RL
Now back to lifelong RL. As we discussed, a fundamental challenge here is the excessive drifting of the weights $\theta_{t}$, and this can be fixed by enforcing the proximity to a good reference point $\theta_{\mathrm{ref}}$. Different from existing approaches like $L_{2}$ regularization and resetting, parameter-free OCO provides a principled solution to this problem without hyperparameter-tuning. Naming this algorithm as Trac, we present its generic template as Algorithm 1, which calls Algorithm 2 (a one-dimensional scale tuner) as the key subroutine.
From the technical perspective, Trac assembles three techniques in parameter-free OCO: the direction-magnitude decomposition from (Cutkosky & Orabona, 2018), the additive aggregation from (Cutkosky, 2019), and the $\mathrm{erfi}$ potential function from (Zhang et al., 2024b). In addition, we follow (Cutkosky et al., 2023) for the recommended default parameter setup.
@@
  • To fix this issue, Trac uses the tuner (Algorithm 2) to select the scaling parameter $S_{t+1}$, making it data-dependent. Typically $S_{t+1}$ is within $[0,1]$ (see Figure 13 to 15), therefore essentially, we define the updated weight $\theta_{t+1}$ as a convex combination of the Base’s weight $\theta_{t}^{\mathrm{Base}}$ and the reference point $\theta_{\mathrm{ref}}$,
@@ 4 Experiment
OpenAI Procgen
We first evaluate on OpenAI Procgen, a suite of 16 procedurally generated game environments (Cobbe et al., 2020). We introduce distribution shifts by sampling a new procedurally generated level of the current game every 2 million time steps, treating each level as a distinct task.
Arcade Learning Environment (ALE) Atari
The Arcade Learning Environment (ALE) Atari 2600 benchmark is a collection of classic arcade games designed to assess reinforcement learning agents’ performance across a range of diverse gaming scenarios (Bellemare et al., 2013). We introduce distribution shifts by switching to a new Atari game every 4 million timesteps, where each game switch introduces a new task. This benchmark is more challenging compared to OpenAI Procgen: it requires the agent to handle distribution shifts in both the input (state) and the target (reward).
Gym Control
We use the CartPole-v1 and Acrobot-v1 environments from the Gym Classic Control suite, along with LunarLander-v2 from Box2d Control. To introduce distribution shifts, a method outlined by Mendez et al. (2020) involves periodically altering the environment dynamics. Although such distribution shifts pose only mild challenges for robust methods like PPO with Adam (Appendix C). We instead implement a more challenging form of distribution shift. Every 200 steps we perturb each observation dimension with random noise within a range of ±$\pm 2$, treating each perturbation phase as a distinct task.
@@ 5 Discussion
Related work Combating loss of plasticity has been studied extensively in lifelong RL. A typical challenge for existing solutions is the tuning of their hyperparameters, which requires prior knowledge on the nature of the distribution shift, e.g., (Asadi et al., 2024; Ben-Iwhiwhu et al., 2022; Nikishin et al., 2024; Sokar et al., 2023; Mesbahi et al., 2024). An architectural modification called CReLU is studied in (Abbas et al., 2023), but our experiments suggest that its benefit might be specific to the Atari setup. Besides, Abel et al. (2018a, b) presented a theoretical analysis of skill transfer in lifelong RL, based on value iteration. Moreover, related contributions in nonstationary RL, where reward and state transition functions also change unpredictably, are limited to theoretical sequential decision-making settings with a focus on establishing complexity bounds (Roy et al., 2019; Cheung et al., 2020; Wei & Luo, 2021; Mao et al., 2020).
@@
Particularly, we are inspired by the Mechanic algorithm from (Cutkosky et al., 2023), which goes beyond the traditional convex setting of parameter-free OCO to handle stationary deep learning optimization tasks. Lifelong reinforcement learning, however, introduces a layer of complexity with its inherent nonstationarity. Furthermore, compared to Mechanic, Trac improves the scale tuner there (which is based on the coin-betting framework; Orabona & Pál, 2016) by the $\mathrm{erfi}$ algorithm that enjoys a better OCO performance guarantee. As an ablation study, we empirically compare Trac and Mechanic in the Appendix E (Figure 16). We find that Trac is slightly better, but both algorithms can mitigate the loss of plasticity, suggesting the effectiveness of the general “parameter-free” principle in lifelong RL.
Trac encourages positive transfer
In our experiments, we observe that Trac’s reward decline due to distribution shifts is less severe than that of baseline methods. These results may suggest that Trac facilitates positive transfer between related tasks in a lifelong learning setting. To investigate this further, we compared Trac to a privileged weight-resetting approach, where the network’s parameters are reset for each new task, in the Gym Control environments (see Appendix A). Our results show that Trac maintains higher rewards during tasks than privileged weight-resetting and avoids declining to the same low reward levels as privileged weight-resetting at the start of a new task (Figure 8).
On the choice of $\theta_{\mathrm{ref}}$
In general, the reference point $\theta_{\text{ref}}$ should be good or “safe” for Trac to perform effectively. One might presume that achieving this requires “warmstarting”, or pre-training using the underlying Base optimizer. While our experiments validate that such warmstarting is indeed beneficial (Appendix B), our main experiments show that even a random initialization of the policy’s weight serves as a good enough $\theta_{\text{ref}}$, even when tasks are similar (Figure 4).
@@
Tuning $L_{2}$ regularization
The success of Trac suggests that an adaptive form of regularization—anchoring to the reference point $\theta_{\mathrm{ref}}$—may suffice to counteract both mild and extreme forms of loss of plasticity. From this angle, we further elaborate the limitation of the $L_{2}$ regularization approach considered in (Kumar et al., 2023). It requires selecting a regularization strength parameter $\lambda$ through cross-validation, which is incompatible with the one-shot nature of lifelong learning settings. Furthermore, it is nontrivial to select the search grid: for example, we tried the $\lambda$-grid suggested by (Kumar et al., 2023), and there is no effective $\lambda$ value within the grid for the lifelong RL environments we consider. All the values are too small.
@@
Figure 7: For each Gym Control environment and the initial ten tasks, we identified the best $\lambda$, which is the regularization strength that maximizes reward for each task’s specific distribution shift. Additionally we determined the best overall (well-tuned) $\lambda$ for each environment, optimizing performance across all tasks (CartPole $\lambda=25$, Acrobot $\lambda=15$, LunarLander $\lambda=45$). The results demonstrate that each environment and each task’s distribution shift is sensitive to different $\lambda$ and that Trac PPO performs competitively with each environment’s well-tuned regularization strength.
Near convexity of lifelong RL
Our results demonstrate the rapid adaptation of Trac, in lifelong RL problems with complicated function approximation. From the perspective of optimization, the latter requires tackling both nonconvexity and nonstationarity, which is typically regarded intractable in theory. Perhaps surprisingly, when approaching this complex problem using the theoretical insights from OCO, we observe compelling results. This suggests a certain “hidden convexity” in this problem, which could be an exciting direction for both theoretical and empirical research (e.g., policy gradient methods provably converge to global optimizers in linear quadratic control (Hu et al., 2023)).
Limitations
While Trac offers robust adaptability in nonstationary environments, it can exhibit suboptimal performance at the outset due to the time it requires to “warm up”. In the early stages of deployment, Trac might underperform compared to the baseline optimizer. We address this by proposing a warmstarting solution detailed in Appendix B, which helps bridge the initial performance gap. Currently, our results demonstrate the effectiveness of Trac in conjunction with PPO. However, Trac is fundamentally an optimizer that could broadly apply to other RL algorithms. Future work should extend the application of Trac to additional algorithms such as Soft Actor-Critic (Haarnoja et al., 2018) and Deep Q-Networks (Mnih et al., 2015).
@@ Appendix B Warmstarting
Appendix C Gravity Based Distribution Shifts

### v3

@@ 1 Introduction
Spot’s struggle exemplifies a well-known and extensively studied challenge in real-world decision making: lifelong reinforcement learning (lifelong RL) Abel et al. (2024); Nath et al. (2023); Mendez et al. (2020); Xie & Finn (2022). In lifelong RL, the learning agent must continually acquire new knowledge to adapt to the nonstationarity of the environment. At first glance, there appears to be an obvious solution: given a policy gradient oracle, the agent could just keep running gradient descent nonstop. However, recent experiments have demonstrated an intriguing behavior called loss of plasticity (Dohare et al., 2021; Lyle et al., 2022; Abbas et al., 2023; Sokar et al., 2023): despite persistent gradient steps, such an agent can gradually lose its responsiveness to incoming observations. There are even extreme cases of loss of plasticity (known as negative transfer or primacy bias), where prior learning can significantly hamper the performance in new tasks (Nikishin et al., 2022; Ahn et al., 2024); see Figure 1 for an example. All these suggest that the problem is more involved than one might think.
@@
  • Algorithm: Trac Building on a series of results in OCO (Cutkosky & Orabona, 2018; Cutkosky, 2019; Cutkosky et al., 2023; Zhang et al., 2024b), we propose a (hyper)-parameter-free optimizer for lifelong RL, called Trac (AdapTive RegularizAtion in Continual environments). Intuitively, the idea is a refinement of regularization: instead of manually selecting the magnitude of regularization beforehand, Trac chooses that in an online, data-dependent manner. From the perspective of OCO theory, Trac is insensitive to its own hyperparameter, which means that no hyperparameter tuning is necessary in practice. Furthermore, as an optimization approach to lifelong RL, Trac is compatible with any policy parameterization method.
  • Experiment Using Proximal Policy Optimization (PPO) (Schulman et al., 2017), we conduct comprehensive experiments on the instantiation of Trac called Trac PPO. A diverse range of lifelong RL environments are tested (based on Procgen, Atari, and Gym Control), with considerably larger scale than prior works. In settings where existing approaches (Abbas et al., 2023; Kumar et al., 2023; Nath et al., 2023) struggle, we find that Trac PPO
@@ 2 Lifelong RL
Deep RL approaches, including PPO (Schulman et al., 2017) and others, crucially utilize the idea of policy parameterization. Specifically, a policy refers to the distribution of the agent’s action $a_{t}$ (conditioned on the historical observations), and we use $\theta_{t}\in\mathbb{R}^{d}$ to denote the parameterizing weight vector. After sampling $a_{t}$ and receiving new observations, the agent could define a loss function $J_{t}(\theta)$ that characterizes the “hypothetical performance” of each weight $\theta\in\mathbb{R}^{d}$. Then, by computing the policy gradient $g_{t}=\nabla J_{t}(\theta_{t})$, one could apply a first order optimization algorithm [^3] Formally, a dynamical system that given its state $\theta_{t}$ and input $g_{t}$ outputs the new state $\textsc{OPT}(\theta_{t},g_{t})$. OPT to obtain the updated weight, $\theta_{t+1}=\textsc{OPT}(\theta_{t},g_{t})$.
@@
In the RL literature, the use of “lifelong” and “continual” varies significantly across studies, which may lead to confusion. Abel et al. (2024) characterized continual reinforcement learning (CRL) as a never-ending learning process. However, much of the literature cited under CRL, such as (Abbas et al., 2023; Ahn et al., 2024), primarily focuses on the problem of backward transfer (avoiding catastrophic forgetting). Various policy-based architectures, such as those proposed by Rolnick et al. (2019); Schwarz et al. (2018); Nath et al. (2023), focus on tackling this issue. Conversely, the present work addresses the problem of forward transfer, which refers to the rapid adaptation to new tasks. Because of this we use “lifelong” rather than “continual” in our exposition, similar to (Thrun, 1996; Abel et al., 2018b; Julian et al., 2020).
@@ 3 Method
TRAC for Lifelong RL: In lifelong RL, a key issue is the excessive drifting of weights $\theta_{t}$, which can detrimentally affect adapting to new tasks. To address this, TRAC enforces proximity to a well-chosen reference point $\theta_{\mathrm{ref}}$, providing a principled solution derived from a decade of research in parameter-free OCO. Unlike traditional methods such as $L_{2}$ regularization or resetting, TRAC avoids hyperparameter tuning, utilizing the properties of OCO to maintain weight stability and manage the drift effectively.
The core of TRAC, similar to other parameter-free optimizers, incorporates three techniques:
  • Direction-Magnitude Decomposition: Inspired by Cutkosky & Orabona (2018), this technique employs a carefully designed one-dimensional algorithm, the "parameter-free tuner," atop a base optimizer. This setup acts as a data-dependent regularizer, controlling the extent to which the iterates deviate from their initialization, thereby minimizing loss of plasticity, which is crucial given the high plasticity at the initial policy parameterization (Abbas et al., 2023).
  • Erfi Potential Function: Building on the previous concept, the tuner utilizes the Erfi potential function, as developed by Zhang et al. (2024a). This function is crafted to effectively balance the distance of the iterates from both the origin and the empirical optimum. It manages the update magnitude by focusing on the gradient projection along the direction $\theta_{t}-\theta_{\mathrm{ref}}$.
  • Additive Aggregation: The tuner above necessitates discounting. Thus, we employ Additive Aggregation by Cutkosky (2019). This approach enables the combination of multiple parameter-free OCO algorithms, each with different discount factors, to approximate the performance of the best-performing algorithm. Importantly, it facilitates the automatic selection of the optimal discount factor during training.
These three components crucially work together to guarantee good regret bounds in the convex setting and are the minimum requirement for any reasonable parameter-free optimizer.
@@
  • To fix this issue, Trac uses the tuner (Algorithm 2) to select the scaling parameter $S_{t+1}$, making it data-dependent. Typically $S_{t+1}$ is within $[0,1]$ (see Figure 17 to 19), therefore essentially, we define the updated weight $\theta_{t+1}$ as a convex combination of the Base’s weight $\theta_{t}^{\mathrm{Base}}$ and the reference point $\theta_{\mathrm{ref}}$,
@@ 4 Experiment
Procgen We first evaluate on OpenAI Procgen, a suite of 16 procedurally generated game environments (Cobbe et al., 2020). We introduce distribution shifts by sampling a new procedurally generated level of the current game every 2 million time steps, treating each level as a distinct task.
One key advantage of Trac is that it functions as an optimizer, making it orthogonal to various policy methods such as PPO, as well as other baselines like Online EWC (Schwarz et al., 2018), IMPALA (Espeholt et al., 2018), Modulating Masks (Nath et al., 2023), and CLEAR (Rolnick et al., 2019). In Appendix C, we evaluate these methods using both Trac and Adam on the Procgen setup. We find that in every environment, Trac improves the performance of these algorithms.
Atari The Arcade Learning Environment (ALE) Atari 2600 benchmark is a collection of classic arcade games designed to assess reinforcement learning agents’ performance across a range of diverse gaming scenarios (Bellemare et al., 2013). We introduce distribution shifts by switching to a new Atari game every 4 million timesteps, where each game switch introduces a new task. This benchmark is more challenging compared to OpenAI Procgen: it requires the agent to handle distribution shifts in both the input (state) and the target (reward).
Gym Control We use the CartPole-v1 and Acrobot-v1 environments from the Gym Classic Control suite, along with LunarLander-v2 from Box2d Control. To introduce distribution shifts, Mendez et al. (2020) periodically alters the environment dynamics. Although such distribution shifts pose only mild challenges for robust methods like PPO with Adam (Appendix D). We instead implement a more challenging form of distribution shift. Every 200 steps we perturb each observation dimension with random noise within a range of ±$\pm 2$, treating each perturbation phase as a distinct task.
@@ 5 Discussion
Related work Combating loss of plasticity has been studied extensively in lifelong RL. A typical challenge for existing solutions is the tuning of their hyperparameters, which requires prior knowledge on the nature of the distribution shift, e.g., (Asadi et al., 2024; Nath et al., 2023; Nikishin et al., 2024; Sokar et al., 2023; Mesbahi et al., 2024). An architectural modification called CReLU is studied in (Abbas et al., 2023), but our experiments suggest that its benefit might be specific to the Atari setup. Besides, Abel et al. (2018a, b) presented a theoretical analysis of skill transfer in lifelong RL, based on value iteration. Moreover, related contributions in nonstationary RL, where reward and state transition functions also change unpredictably, are limited to theoretical sequential decision-making settings with a focus on establishing complexity bounds (Roy et al., 2019; Cheung et al., 2020; Wei & Luo, 2021; Mao et al., 2020).
@@
Particularly, we are inspired by the Mechanic algorithm from (Cutkosky et al., 2023), which goes beyond the traditional convex setting of parameter-free OCO to handle stationary deep learning optimization tasks. Lifelong reinforcement learning, however, introduces a layer of complexity with its inherent nonstationarity. Furthermore, compared to Mechanic, Trac improves the scale tuner there (which is based on the coin-betting framework; Orabona & Pál, 2016) by the $\mathrm{erfi}$ algorithm that enjoys a better OCO performance guarantee. As an ablation study, we empirically compare Trac and Mechanic in the Appendix G (Table 3). We find that Trac is slightly better, but both algorithms can mitigate the loss of plasticity, suggesting the effectiveness of the general “parameter-free” principle in lifelong RL.
Trac encourages positive transfer In our experiments, we observe that Trac’s reward decline due to distribution shifts is less severe than that of baseline methods. These results may suggest Trac facilitates positive transfer between related tasks. To investigate this further, we compared Trac to a privileged weight-resetting approach, where the network’s parameters are reset for each new task, in the Gym Control environments (see Appendix A). Our results show that Trac maintains higher rewards during tasks than privileged weight-resetting and avoids declining to the same low reward levels as privileged weight-resetting at the start of a new task (Figure 8).
On the choice of $\theta_{\mathrm{ref}}$ In general, the reference point $\theta_{\text{ref}}$ should be good or “safe” for Trac to perform effectively. One might presume that achieving this requires “warmstarting”, or pre-training using the underlying Base optimizer. While our experiments validate that such warmstarting is indeed beneficial (Appendix B), our main experiments show that even a random initialization of the policy’s weight serves as a good enough $\theta_{\text{ref}}$, even when tasks are similar (Figure 4).
@@
Tuning $L_{2}$ regularization The success of Trac suggests that an adaptive form of regularization—anchoring to the reference point $\theta_{\mathrm{ref}}$—may suffice to counteract both mild and extreme forms of loss of plasticity. From this angle, we further elaborate the limitation of the $L_{2}$ regularization approach considered in (Kumar et al., 2023). It requires selecting a regularization strength parameter $\lambda$ through cross-validation, which is incompatible with the one-shot nature of lifelong learning settings. Furthermore, it is nontrivial to select the search grid: for example, we tried the $\lambda$-grid suggested by (Kumar et al., 2023), and there is no effective $\lambda$ value within the grid for the lifelong RL environments we consider. All the values are too small.
@@
Trac compared to other plasticity methods Both layer normalization and plasticity injection Nikishin et al. (2024); Lyle et al. (2023) have been shown to combat plasticity loss. For instance, Appendix E Figure 15 demonstrates that both layer normalization and plasticity injection are effective at reducing plasticity loss when applied to the CartPole environment using Adam as a baseline optimizer. We implemented plasticity injection following the methodology laid out by Nikishin et al. (2024), where plasticity is injected at the start of every distribution shift. While this approach does help in reducing the decline in performance due to plasticity loss, our results indicate that it is consistently outperformed by TRAC across all three control environments—CartPole, Acrobot, and LunarLander. Moreover, while layer normalization improves Adam’s performance, it too is outperformed by TRAC across the same control settings (Figure 15). Notably, combining layer normalization with TRAC resulted in the best performance gains.
Figure 7: For each Gym Control environment and the initial ten tasks, we identified the best $\lambda$, which is the regularization strength that maximizes reward for each task’s specific distribution shift. We also determined the best overall (well-tuned) $\lambda$ for each environment. The results demonstrate that each environment and each task’s distribution shift is sensitive to different $\lambda$ and that Trac PPO performs competitively with each environment’s well-tuned $\lambda$.
Near convexity of lifelong RL Our results demonstrate the rapid adaptation of Trac, in lifelong RL problems with complicated function approximation. From the perspective of optimization, the latter requires tackling both nonconvexity and nonstationarity, which is typically regarded intractable in theory. Perhaps surprisingly, when approaching this complex problem using the theoretical insights from OCO, we observe compelling results. This suggests a certain “hidden convexity” in this problem, which could be an exciting direction for both theoretical and empirical research (e.g., policy gradient methods provably converge to global optimizers in linear quadratic control (Hu et al., 2023)).
Limitations While Trac offers robust adaptability in nonstationary environments, it can exhibit suboptimal performance at the outset. In the early stages of deployment, Trac might underperform compared to the baseline optimizer. We address this by proposing a warmstarting solution detailed in Appendix B, which helps increase the initial performance gap.
@@ Appendix B Warmstarting
Appendix C Other RL Baselines
While PPO is a widely used policy gradient method in reinforcement learning, it is not the only approach applicable to lifelong RL. Other continual RL methods, such as IMPALA (Espeholt et al., 2018), Online EWC (Schwarz et al., 2018), CLEAR (Rolnick et al., 2019), and Modulating Masks (Nath et al., 2023), are designed to address challenges like catastrophic forgetting in dynamic, nonstationary environments. We incorporated these algorithm implementations adapted from the code from Nath et al. (2023) into our experiments to offer a more comprehensive evaluation. These methods vary in their mechanisms for maintaining task performance over time but may still suffer from plasticity loss in later stages of training.
Mitigating plasticity loss across policy methods: Figure 12 demonstrates the performance of various continual RL methods when paired with Adam and Trac optimizers. The results indicate that when using Adam, methods like IMPALA, Online EWC, CLEAR, and Modulating Masks exhibit a noticeable decline in performance over time due to plasticity loss, particularly in later levels of the Procgen environments. In contrast, pairing these methods with Trac instead of Adam leads to significant improvements, mitigating plasticity loss and enhancing reward performance across subsequent distribution shifts.
To quantify these improvements, Figures 13, 12 present the average normalized rewards over five seeds and 120M timesteps for each method across four different Procgen environments: Starpilot, Dodgeball, Chaser, and Fruitbot. Across all environments, methods that use Trac outperform their Adam-based counterparts, consistently maintaining higher rewards over time.
Figure 12: Performance comparison between Adam-based and Trac-based continual RL methods (IMPALA, Online EWC, CLEAR, Modulating Masks) in Starpilot. While Adam suffers from plasticity loss in later levels, Trac effectively mitigates this and maintains better performance over distribution shifts. For clarity, standard deviation fills are omitted here but included in the bar plot.
Figure 13: Average normalized rewards over five seeds and 120M timesteps for Dodeball, Chaser, and Fruitbot. Each method (IMPALA, Online EWC, CLEAR, and Modulating Masks) is evaluated using both Adam and Trac. Trac consistently outperforms Adam across all methods and environments.
On average, across the Procgen environments, Trac led to performance improvements over Adam by the following margins: 21.83% for IMPALA, 15.86% for Online EWC, 14.41% for CLEAR, and 10.14% for Modulating Masks.
General Applicability of Trac: It is important to highlight that Trac is orthogonal to the learning or policy algorithms themselves. It can be seamlessly integrated into various reinforcement learning architectures by simply replacing their optimizer (e.g., Adam or RMSPROP). Our results demonstrate that Trac enhances performance across different algorithms and environments, consistently outperforming Adam in mitigating plasticity loss.
Appendix D Gravity Based Distribution Shifts